OpenStudy (anonymous):

How much interest will you earn if you invest $10,000 at 7% for 2 years? Use the formula: I=prt Select one: a. $1200 b. $1300 c. $1400 d. $1500

OpenStudy (haseeb96):

C is the correct answer. I=0.07×10,000× 2 I=1400

OpenStudy (anonymous):

Could You Explain How Please?

OpenStudy (haseeb96):

P showing percentage so here in the question 7%= 7/100 R showing money T showing time so that is given in years which is 22 now the formula of interest is I=prt now just put the values in the formula and get the answer.

OpenStudy (haseeb96):

Sorry 2 years is given.

OpenStudy (anonymous):

Alright Thanks